OpenStudy (anonymous):

What is true about the solutions of a quadratic equation when the radicand of the quadratic formula is zero? No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions

OpenStudy (amistre64):

lets give it a shot:\[x=\frac{-b\pm\sqrt{0}}{2a}\]

OpenStudy (amistre64):

-b+0 = -b -b-0 = -b so that simplifies to what ...\[x=\frac{-b}{2a}\]

OpenStudy (anonymous):

two identical solutions (x = -b/2a)

OpenStudy (anonymous):

thnk u

OpenStudy (amistre64):

i was hoping they could come to that realization on their own :)